OpenStudy (anonymous):

Two trains, one 350 feet long, the other 450 feet long, on parallel tracks, can pass each other completely in 8 seconds when moving in opposite directions. When moving in the same direction, the faster train completely passes the slower one in 16 seconds. Find the speed of the slower train.

OpenStudy (barrycarter):

I like this problem. How far have you gotten?

OpenStudy (anonymous):

i haven't a clue on where to start.

OpenStudy (barrycarter):

OK, this will work better on a whiteboard. Can you meet me at: http://www.scribblar.com/wwcdp0qj

OpenStudy (kropot72):

The combined speed of the trains in feet per second is: \[\frac{350+450}{8}\] The extra speed of the faster train in feet per second is: \[\frac{350+450}{16}\] Let the speed of the slower train = x feet per second. Then \[x+(x+\frac{350+450}{16})=\frac{350+450}{8}\] which simplifies to: \[2x+\frac{800}{16}=\frac{800}{8}\] Can you now solve for x (the speed of the slower train in feet per second)?